OpenStudy (anonymous):

What rate is equal to 1 mile/8 minutes

OpenStudy (raden):

do u want converting it to mil/hour ?

OpenStudy (anonymous):

yup

OpenStudy (raden):

okay, it can be 1 mile/8 minutes = 1 mile/(8/60) h = 1 * 60/8 mile/h = 30/4 = 7.5 mile/h

OpenStudy (anonymous):

thnks

OpenStudy (raden):

yw